# Different Inhomogeneous Evolutionary Histories for Uranus and Neptune

### v1

@@ Abstract
We present updated non-adiabatic and inhomogeneous evolution models for Uranus and Neptune, employing an interior composition of methane, ammonia, water, and rocks. Following formation trends of the gas giants, Uranus and Neptune formation models are applied, where both planets begin with layers stable to convection. Both planets are subject to convective mixing throughout their evolution. Uranus undergoes modest convective mixing, preserving much of its primordial internal heat. In contrast, Neptune’s interior undergoes extensive mixing, homogenization, and adiabatic cooling of the outer 40% of its envelope. The subsequent release of internal energy in Neptune, driven by the convective instability of its primordial outer compositional gradient, accounts for its higher luminosity relative to Uranus. Thus, the observed luminosity differences between Uranus and Neptune could be primarily dictated by the convective stability of their outer envelopes. The extensive convective mixing in Neptune leads to a higher metallicity in its outer region compared to Uranus, a feature seen in atmospheric measurements and shown in past interior models of Neptune. Due to Neptune’s more pronounced cooling, our models predict favorable conditions for hydrogen-water immiscibility in its envelope.
@@ 1 Introduction
Uranus and Neptune remain the least understood gas giant planets in our Solar System. Unlike Jupiter and Saturn, extensively studied by spacecraft such as Juno (Bolton et al., 2017) and Cassini (Matson et al., 2003), Uranus and Neptune have been visited only once by Voyager 2, in 1986 and 1989, respectively. Despite the limited observational data, valuable insights have been gained into their internal structures of Uranus and Neptune. Over the past three decades, exoplanet research has uncovered that most of the exoplanet masses and sizes in our galaxy are between the size of the Earth and Neptune (Winn & Fabrycky, 2015; Dressing & Charbonneau, 2015; Christiansen et al., 2023; Parc et al., 2024). Hence, understanding their formation, interior structure, and evolution is becoming evermore relevant.
The difference between the luminosities of Uranus and Neptune is one of the most significant questions regarding both planets. Early studies indicated that Neptune’s luminosity could be accounted for using relatively simple adiabatic models (Hubbard, 1978). At the same time, similar early three-layer models for Uranus predicted cooling times longer than those implied by its measured effective temperature (Hubbard, 1978; Hubbard et al., 1995; Fortney et al., 2011; Scheibe et al., 2019). Consequently, the near-zero internal flux observed in Uranus remained a longstanding puzzle (Pearl & Conrath, 1991; Podolak et al., 1995; Hubbard et al., 1995; Fortney et al., 2011), particularly given the similar mass, radius, atmospheric metallicity, and magnetic field characteristics of the two planets (Guillot & Gautier, 2015). Traditionally, evolutionary models for Uranus and Neptune have employed a three-layer adiabatic structure consisting of a rocky core, a water-rich envelope, and an outer hydrogen-helium (H-He) layer (e.g., Hubbard, 1968, 1977, 1978; Hubbard & Macfarlane, 1980; Nettelmann et al., 2013). To address Uranus’s low luminosity, Nettelmann et al. (2016) introduced a thermal boundary layer (TBL), and positioned it in the outer layers of the planet. This TBL restricts internal heat transport, thereby enabling efficient cooling of the outer envelope, while conserving heat within the deep interior (also see Scheibe et al., 2019, 2021).
@@
Subsequent research suggested that incorporating higher H-He abundances with ammonia-methane-water ice mixtures may further refine interior models of Uranus and Neptune (Movshovitz & Fortney, 2022; Lin et al., 2024). Particularly, Lin et al. (2024) found a stronger consistency of ice-H-He mixtures with the gravity data ($J_{2},\ J_{4}$) of Uranus. Independently, formation models by Valletta & Helled (2022, VH22) also predicted significant mixing of ices with H-He, suggesting H-He fractions between 20% and 40% in the planetary envelopes of the nacent Uranus and Neptune.
@@
In this study, we present updated evolutionary models of Uranus and Neptune guided by formation scenarios proposed by VH22. We aim here to explain the luminosity differences between these ice giants via evolutionary processes. Previous hypotheses for the low intrinsic luminosity of Uranus included silicates in Uranus’s atmosphere (Hubbard & Macfarlane, 1980), the presence of a freezing core (Stixrude et al., 2021), and enhanced conduction (see discussion in Scheibe et al., 2021). Yet, none have concurrently addressed the history of the radius, effective temperature, luminosity, and gravitational harmonics consistently. While past studies have focused primarily on Uranus, our research here investigates how convective mixing impacts luminosity and effective and internal temperature evolutionary trajectories of both planets, incorporating ammonia-methane-water and rock mixtures throughout their envelopes.
@@
Note. — Gravity data for Uranus and Neptune come from Jacobson (2025) and Jacobson (2009), respectively. The effective temperature and Bond albedo measurements are derived by Pearl & Conrath (1991, PC91) from Voyager 2 data. Likewise, the emitted power, absorbed power, and intrinsic flux was measured by PC91. Recently, Wang et al. (2025, W25) measured a higher intrinsic flux from Uranus based on seasonal variations of the planet. A new Bond albedo was inferred by Irwin et al. (I25 2025), who also found a higher intrinsic flux for Uranus. The rotation periods of both planets are tabulated as measured by Voyager 2 (V2; Desch et al., 1986; Warwick et al., 1986, 1989) and as revised by Helled et al. (2010, H10).
@@ 2.1 Equations of State
We use the hydrogen-helium (H-He) equation of state (EOS) from Chabrier & Debras (2021), which is based on the EOS of Chabrier et al. (2019) and incorporates the non-ideal entropy and density effects of Militzer & Hubbard (2013). All models in this work have a constant $Y/(X+Y)=0.277$, consistent with solar abundance ratios (Bahcall et al., 2006), where $X$ and $Y$ are the relative hydrogen and helium mass fractions, . In contrast to previous pioneering studies that modeled Uranus and Neptune interiors exclusively with pure water and rocks (Scheibe et al., 2019, 2021; Vazan & Helled, 2020), we adopt a solar composition mixture comprising methane, ammonia, and water in a 4:1:7 ratio (Asplund et al., 2009; Bethkenhagen et al., 2017) using the volume-addition law (See discussions in Saumon et al., 1995; Tejada Arevalo et al., 2024). Bethkenhagen et al. (2017) found deviations of the volume-addition law up to 4% from direct mixture EOS calculations, justifying its use here. For the ice mixture, we use the multi-phase water EOS provided by Haldemann et al. (2020), equipped with the updated water EOS corrections from Mazevet et al. (2019). The ammonia and methane EOSes are those calculated by Bethkenhagen et al. (2017). We transition smoothly to the analytic EOS fits computed by Gao et al. (2023) for ammonia and Setzmann & Wagner (1991) for methane to capture lower temperature and density conditions (below the lower limits of 1000 K and 0.6 g cm-3; Bethkenhagen et al., 2017). [^1] The zero points of the Gao et al. (2023) and Setzmann & Wagner (1991) EOSes are chosen to be consistent with the zero point of the EOSes of Bethkenhagen et al. (2017).
@@
Figure 1 illustrates the resulting temperature and density differences between H-He-water and H-He-ice mixtures ($Z=0.8$) at various specific entropies. The isentropic differences in temperature and density are shown in Figure 1 between H-He-water and H-He-ice mixtures are several $10^{3}$ K and $\gtrsim 5\%$ in density. These differences are significant enough to warrant the use of ice mixtures instead of pure water. We, hence, argue here that a pure water EOS does not sufficiently approximate ice mixtures to justify using pure water in Uranus and Neptune models. For rocky materials, we use the MgSiO3 EOS originally derived by Keane (1954), including the vibrational and thermal effects as calculated and applied by Zhang & Rogers (2022). We omit considerations of iron content, since current evidence suggests that the ice giants likely possess minimal iron in their cores or envelopes (Movshovitz & Fortney, 2022; Malamud et al., 2024). A more detailed discussion of the H-He-Z mixture EOS can be found in Tejada Arevalo et al. (2024).
@@ 2.2 Atmospheres
of which $A$ is the Bond albedo, $d$ is the semi-major axis, and $T_{\odot}$ and $R_{\odot}$ are the effective temperature and radius of the Sun respectively. Throughout this work we apply the coefficients $K_{U}=1.481$ and $K_{N}=1.451$ in Equation 1, Bond albedos of 0.3 and 0.29 (Pearl & Conrath, 1991), and $d=19.19$ A.U. and $d=30.07$ A.U. in Equation 3 for Uranus and Neptune respectively, as done in Scheibe et al. (2021). The intrinsic luminosity flux of the evolution models is hence calculated via the Stefan-Boltzmann law:
@@ 2.3 Evolution and Transport
The conductive heat transport implemented in APPLE is comprehensively described in Sur et al. (2024). In this study, we adopt the multi-phase thermal conductivity model of French & Redmer (2017) for the heavy-element thermal conduction component. Besides thermal conduction, compositional diffusion is calculated via the diffusion equation is detailed in Section 5 of Sur et al. (2024).
@@ 3 Evolution Models With Varying Entropy Profiles
Figure 3 illustrates the effects of varying interior specific entropy values from 3 to 3.7 kB baryon-1 in Uranus-mass models, analogous to the approach of Section 3.2 in Tejada Arevalo et al. (2025) for Jupiter. All models in this Section are initialized at the same outer entropy of 5.5 kB baryon-1. The differences in initial interior entropy profiles lead to distinct evolutionary pathways in temperature and composition, with higher initial entropies promoting extensive convective mixing. The bottom panel of Figure 3 displays the radius, effective temperature, luminosity, outer heavy-element abundance, and gravity harmonic evolution of each model. Higher luminosities in these models are attributed to rapid heat release from zones interior to the outer compositional gradient into the external layers, particularly evident in the hottest model (red curve), which undergoes near-total convective mixing. The hottest model undergoes a sudden convective mixing event which causes a dramatic, yet temporary, increase in radius, effective temperature, and intrinsic flux. The heat interior to its initial outer stabilizing $Z$ gradient is released upon mixing when the model reaches $\sim$0.4 Gyr, at which the outer regions ($\gtrsim 0.6$ MP, where MP is the planet mass) begin to cool adiabatically. Adiabatic cooling of the outer regions leads to greater internal flux and significantly smaller radii, as shown in the bottom center panel of Figure 3. In contrast, maintaining the outer compositional gradient results in significantly lower luminosities and larger radii. These example models serve illustrative purposes only, highlighting the impacts of varied initial entropy gradients rather than attempting to precisely reproduce Uranus’s observed properties.
@@
Figure 3: Evolution of three Uranus-mass models with initial interior entropies ($S_{i}$) of 3, 3.5, and 3.7 kB baryon-1. The top row shows the initial and present-age profiles of the entropy (left), heavy element abundance (center), and temperature (right). The bottom row shows the evolution of the effective temperature and intrinsic flux (left), radius, and outer $Z$ (center), and $J_{2}$, $J_{4}$ (left). Higher internal entropy profiles yield higher degrees of convective mixing. The blue and green models preserve an outer $Z$ profile gradient, keeping interior temperatures high, displaying larger radii, leading to low outer intrinsic fluxes. The red model is the hottest and does not preserve an outer composition gradient. The red model mixes almost entirely by 0.4 Gyr. This mixing event causes an ephemeral yet dramatic increase in effective temperature, luminosity, and radius. After the mixing of the outer regions, the outer 40% of the structure cools adiabatically since it becomes homogeneous and fully convective. The heat escape subsequent to convective mixing yields smaller radii (bottom center) and higher final intrinsic fluxes (bottom left). The evolution of the outer heavy element abundance of the red model increases from the initial 25% to 78% by mass during the mixing event. The blue and green models evolve from the initial 25% to 30% and 33% by mass fraction. The evolution of $J_{2}$ and $J_{4}$ is significantly different from models which do not mix (bottom right). The interior temperatures at 1 Mbar of the red model are $\sim$8,000 K colder than the unmixed models at 4.56 Gyr. Generally, this figure shows that different entropy initial profiles can either preserve or mix composition profiles, leading to order-of-magnitude intrinsic flux differences.

### v3

@@ Abstract
We present updated non-adiabatic and inhomogeneous evolution models for Uranus and Neptune, employing an interior composition of methane, ammonia, water, and rocks. Following formation trends of the gas giants, Uranus and Neptune formation models are applied, where both planets begin with layers stable to convection. Both planets are subject to convective mixing throughout their evolution. Consistent with past work on this subject, the interior heat of Uranus evolution models is preserved by the stability of an outer composition gradient at lower initial entropy, where convective mixing is inhibited over evolutionary timescales. In contrast, if Neptune’s initial entropy is enough to convectively mix its envelope, it undergoes homogenization and adiabatic cooling of the outer 40% of its envelope. The subsequent release of internal energy during Neptune’s evolution, driven by the convective instability of its primordial outer compositional gradient, accounts for its higher luminosity relative to Uranus. This work proposes that the observed luminosity differences between Uranus and Neptune could be explained by the convective stability of their outer envelopes. The extensive convective mixing in Neptune can lead to a higher metallicity in its outer region compared to Uranus, a feature seen in atmospheric measurements and shown in past interior models of Neptune. Due to Neptune’s more pronounced cooling, our models predict favorable conditions for hydrogen-water immiscibility in its envelope.
@@ 1 Introduction
Uranus and Neptune remain the least understood gas giant planets in our Solar System. Unlike Jupiter and Saturn, extensively studied by spacecraft such as Juno (Bolton et al., 2017) and Cassini (Matson et al., 2003), Uranus and Neptune have been visited only once by Voyager 2, in 1986 and 1989, respectively. Despite the limited observational data, valuable insights have been gained into the internal structures of Uranus and Neptune. Over the past three decades, exoplanet research has uncovered that most of the exoplanet masses and sizes in our galaxy are between the size of the Earth and Neptune (Winn & Fabrycky, 2015; Dressing & Charbonneau, 2015; Christiansen et al., 2023; Parc et al., 2024), with recent exoplanet population studies finding that Neptune-sized objects are among the most common (Fulton & Petigura, 2018) and can serve as calibrations for exoplanet observations (Hasler et al., 2024). Hence, understanding their formation, interior structure, and evolution is becoming ever more relevant.
The difference between the luminosities of Uranus and Neptune is one of the most significant questions regarding both planets. Early studies indicated that Neptune’s luminosity could be accounted for using relatively simple adiabatic models (Hubbard, 1978). At the same time, similar early three-layer models for Uranus predicted cooling times longer than those implied by its measured effective temperature (Hubbard, 1978; Hubbard et al., 1995; Fortney et al., 2011; Scheibe et al., 2019). Consequently, the near-zero internal flux observed in Uranus remained a longstanding puzzle (Pearl & Conrath, 1991; Podolak et al., 1995; Hubbard et al., 1995; Fortney et al., 2011), particularly given the similar mass, radius, atmospheric metallicity, and magnetic field characteristics of the two planets (Guillot & Gautier, 2015). Traditionally, evolutionary models for Uranus and Neptune have employed a three-layer adiabatic structure consisting of a rocky core, a water-rich envelope, and an outer hydrogen-helium (H-He) layer (e.g., Hubbard, 1968, 1977, 1978; Hubbard & Macfarlane, 1980; Nettelmann et al., 2013). To address Uranus’s low luminosity, Nettelmann et al. (2016) introduced a thermal boundary layer (TBL), and positioned it in the outer layers of the planet. This TBL restricts internal heat transport, thereby enabling efficient cooling of the outer envelope while conserving heat within the deep interior (also see Scheibe et al., 2019, 2021).
@@
Subsequent research suggested that incorporating higher H-He abundances with ammonia-methane-water ice mixtures may further refine interior models of Uranus and Neptune (Movshovitz & Fortney, 2022; Lin et al., 2024). Particularly, Lin et al. (2024) found a stronger consistency of ice-H-He mixtures with the gravity data ($J_{2},\ J_{4}$) of Uranus. Independently, formation models by Valletta & Helled (2022, VH22) also predicted significant mixing of ices with H-He, suggesting H-He fractions between 20% and 40% in the planetary envelopes of the nascent Uranus and Neptune.
@@
In this study, we present updated evolutionary models of Uranus and Neptune guided by formation scenarios proposed by VH22. We aim here to explain the luminosity differences between these ice giants via evolutionary processes. Previous hypotheses for the low intrinsic luminosity of Uranus included silicates in Uranus’s atmosphere (Hubbard & Macfarlane, 1980), giant impacts (Reinhardt et al., 2020), the presence of a freezing core (Stixrude et al., 2021), water-hydrogen phase separation (Bailey & Stevenson, 2021), and enhanced conduction (see discussion in Scheibe et al., 2021). Yet, none have concurrently addressed the history of the radius, effective temperature, luminosity, and gravitational harmonics consistently. While past studies have focused primarily on Uranus, our research here investigates how convective mixing impacts luminosity and effective and internal temperature evolutionary trajectories of both planets, incorporating ammonia-methane-water and rock mixtures throughout their envelopes.
@@
Note. — Gravity data for Uranus and Neptune come from Jacobson (2025) and Jacobson (2009), respectively. The effective temperature and Bond albedo measurements are derived by Pearl & Conrath (1991, PC91) from Voyager 2 data. Likewise, the emitted power, absorbed power, and intrinsic flux was measured by PC91. Recently, Wang et al. (2025, W25) measured a higher intrinsic flux from Uranus based on seasonal variations of the planet. A new Bond albedo was found by Irwin et al. (I25 2025), who also found a higher intrinsic flux for Uranus. The rotation periods of both planets are tabulated as measured by Voyager 2 (V2; Desch et al., 1986; Warwick et al., 1986, 1989) and as revised by Helled et al. (2010, H10).
@@ 2.1 Equations of State
We use the hydrogen-helium (H-He) equation of state (EOS) from Chabrier & Debras (2021), which is based on the EOS of Chabrier et al. (2019) and incorporates the non-ideal entropy and density effects of Militzer & Hubbard (2013). All models in this work have a constant $Y/(X+Y)=0.277$, consistent with solar abundance ratios (Bahcall et al., 2006), where $X$ and $Y$ are the relative hydrogen and helium mass fractions. In contrast to previous pioneering studies that modeled Uranus and Neptune interiors exclusively with pure water and rocks (Scheibe et al., 2019, 2021; Vazan & Helled, 2020), we adopt a solar composition mixture comprising methane, ammonia, and water in a 4:1:7 ratio (Asplund et al., 2009; Bethkenhagen et al., 2017) using the volume-addition law (See discussions in Saumon et al., 1995; Tejada Arevalo et al., 2024). Bethkenhagen et al. (2017) found deviations of the volume-addition law up to 4% from direct mixture EOS calculations, justifying its use here. The impact of different ice ratios is discussed and shown in Appendix A. Recently, evolution models of Uranus and Neptune incorporated H-He and pure water mixtures (Scheibe et al., 2019, 2021) and rocks (Vazan & Helled, 2020). The motivation to use ice mixtures instead of pure water comes from the expectation that ice mixtures are expected in both of these planets since their interior conditions are favorable (Bethkenhagen et al., 2017; Cheng et al., 2023; Lin et al., 2024) and their atmospheric metallicity abundances show evidence for ices (Karkoschka, 1994, 1998). For the ice mixture, we use the multi-phase water EOS provided by Haldemann et al. (2020, ;AQUA), equipped with the updated water EOS corrections from Mazevet et al. (2019). [^1] Aguichine et al. (2025) noted that a correction to the entropy values of Mazevet et al. (2019) EOS has been issued by Mazevet et al. (2021). This correction does not affect the density. We calculated models using a version of AQUA replacing the Mazevet et al. (2019) EOS with the Mazevet et al. (2021) EOS, finding no significant changes to the evolution models and no changes to the broader conclusions presented here. See Appendix A, Figure 7, for a discussion and examples. The ammonia and methane EOSes are those calculated by Bethkenhagen et al. (2017). We transition smoothly to the analytic EOS fits computed by Gao et al. (2023) for ammonia and Setzmann & Wagner (1991) for methane to capture lower temperature and density conditions (below the lower limits of 1000 K and 0.6 g cm-3; Bethkenhagen et al., 2017). [^2] The zero points of the Gao et al. (2023) and Setzmann & Wagner (1991) EOSes are chosen to be consistent with the zero point of the EOSes of Bethkenhagen et al. (2017).
@@
Figure 1 illustrates the resulting temperature and density differences between H-He-water and H-He-ice mixtures ($Z=0.8$) at various specific entropies. The isentropic differences in temperature and density shown in Figure 1 between H-He-water and H-He-ice mixtures are several $10^{3}$ K and $\gtrsim 5\%$ in density. These differences are significant enough to warrant the use of ice mixtures instead of pure water. We, hence, argue here that a pure water EOS does not sufficiently approximate ice mixtures to justify using pure water in Uranus and Neptune models. For rocky materials, we use the MgSiO3 EOS originally derived by Keane (1954), including the vibrational and thermal effects as calculated and applied by Zhang & Rogers (2022). We omit considerations of iron content, since current evidence suggests that the ice giants likely possess minimal iron in their cores or envelopes (Movshovitz & Fortney, 2022; Malamud et al., 2024). A more detailed discussion of the H-He-Z mixture EOS can be found in Tejada Arevalo et al. (2024).
@@ 2.2 Atmospheres
of which $A$ is the Bond albedo, $d$ is the semi-major axis, and $T_{\odot}$ and $R_{\odot}$ are the effective temperature and radius of the Sun respectively. Throughout this work, we apply the coefficients $K_{U}=1.481$ and $K_{N}=1.451$ in Equation 1, Bond albedos of 0.3 and 0.29 (Pearl & Conrath, 1991), and $d=19.19$ A.U. and $d=30.07$ A.U. in Equation 3 for Uranus and Neptune, respectively, as done in Scheibe et al. (2021). The intrinsic luminosity flux of the evolution models is hence calculated via the Stefan-Boltzmann law:
@@ 2.3 Evolution and Transport
The conductive heat transport implemented in APPLE is comprehensively described in Sur et al. (2024). In this study, we adopt the multi-phase thermal conductivity model of French (2019) for the heavy-element thermal conduction component. Besides thermal conduction, compositional diffusion is calculated via the diffusion equation is detailed in Section 5 of Sur et al. (2024).
@@ 3 Evolution Models With Varying Entropy Profiles
Figure 3 illustrates the effects of varying interior specific entropy values from 3 to 3.9 kB baryon-1 in Uranus-mass models, analogous to the approach of Section 3.2 in Tejada Arevalo et al. (2025) for Jupiter. Recent evolution studies and static interior work have found that Uranus could harbor interior temperatures of tens of thousands of kelvin provided its interior is non-adiabatic (Vazan & Helled, 2020; Scheibe et al., 2021; Neuenschwander et al., 2024; Lin et al., 2024). These initial entropy ranges correspond to a hot Uranian interior, further justifying our choice of interior entropy. All models in this Section are initialized at the same outer entropy of 5.5 kB baryon-1.444The outermost entropy values have no long-term stabilizing effect, see Knierim & Helled (2024) and Tejada Arevalo et al. (2025) for discussions. The differences in initial interior entropy profiles lead to distinct evolutionary pathways in temperature and composition, with higher initial entropies promoting extensive convective mixing. The bottom panel of Figure 3 displays the radius, effective temperature, luminosity, outer heavy-element abundance, and gravity harmonic evolution of each model. Higher luminosities in these models are attributed to rapid heat release from zones interior to the outer compositional gradient into the external layers, particularly evident in the hottest model (red curve), which undergoes near-total convective mixing. The hottest model undergoes a sudden convective mixing event which causes a dramatic, yet temporary, increase in radius, effective temperature, and intrinsic flux. The heat interior to its initial outer stabilizing $Z$ gradient is released upon mixing when the model reaches $\sim$0.4 Gyr, at which the outer regions ($\gtrsim 0.6$ MP, where MP is the planet mass) begin to cool adiabatically. Adiabatic cooling of the outer regions leads to greater internal flux and significantly smaller radii, as shown in the center and center bottom panels of Figure 3. In contrast, maintaining the outer compositional gradient results in significantly lower luminosities and larger radii. These example models serve illustrative purposes only, highlighting the impacts of varied initial entropy gradients rather than attempting to precisely reproduce Uranus’s observed properties.
The demonstration models in Figure 3 show that a sufficiently high initial interior entropy will mix the envelope, leading to systematically higher luminosities and smaller radii than models which do not mix (also shown in Appendix B, Figure 9). These models also show that to keep a stabilizing outer layer and better match the radius of a Uranus-mass model, its interior entropy (for this ice mixture ratio) should be $\sim$3 kB baryon-1. This corresponds to initial interior temperatures of $\sim$15,000 K. When the heavy constituents are restricted to pure water mixed with H–He, comparable results are obtained with an entropy of roughly 2.4 kB baryon-1 (See Appendix A).
Figure 3 further shows that entropies of $\sim$3.8 kB baryon-1 undermine the stabilizing $Z$ gradient, and larger values do so at progressively earlier ages. Models that preserve the stable layer (green curve) emit higher luminosities than those that do not (red curve), yet they predict noticeably smaller radii, as seen in the bottom-center panel. The final parameters considered in this study are visualized as a function of initial interior entropy in Appendix B, Figure 9. Entropies $\gtrsim$ 3.9 kB baryon-1 fully homogenize the interior. For a pure-water EOS with H–He, destabilization requires about 3 kB baryon-1. Mixtures with lower methane–ammonia content thus need entropies between these bounds, whereas mixtures richer in methane and ammonia demand even higher values to erase the outer compositional gradient.
@@
Figure 3: Evolution of Uranus-mass models with initial interior entropies ($S_{i}$) ranging between 3 and 3.9 kB baryon-1. The top row shows the initial and present-age profiles of the entropy (left), heavy element abundance (center), and temperature (right). The bottom row shows the evolution of the effective temperature and intrinsic flux (left), radius, and outer $Z$ (center), and $J_{2}$, $J_{4}$ (left). The magenta, blue, and green models preserve an outer $Z$ profile gradient, keeping interior temperatures high, displaying larger radii, leading to low outer intrinsic fluxes. The orange and red model are the hottest and do not preserve an outer composition gradient. The red model mixes almost entirely by 0.3 Gyr and the orange by 0.4 Gyr. This mixing event causes an ephemeral yet dramatic increase in effective temperature, luminosity, and radius. Note that while the green model achieve similarly high final luminosities (center left), it evolves to different radii (green is inflated). This is discussed and further visualized in Appendix B. Models at or higher initial interior entropy than $\sim$3.7 convectively mix the outer regions. The heat escape subsequent to convective mixing yields smaller radii (bottom center) and higher final intrinsic fluxes (bottom right). The evolution of the outer heavy element abundance of the red model increases from the initial 25% to 78% by mass during the mixing event (dashed lines, bottom center). The evolution of $J_{2}$ and $J_{4}$ is significantly different from models which do not mix (bottom right).